OpenStudy (anonymous):

Find the derivative of y=sin^2x The answer is sin2x, but I don't get why that is.

OpenStudy (campbell_st):

ok... so let u = sin(x) du/dx = cos(x) so y = u^2 then dy/du = 2u then using the chain rule \[\frac{dy}{dx} = \frac{dy}{du} \times \frac{du}{dx}\] which is \[\frac{dy}{dx} = 2u \times \cos(x)\] make the substitution u = sin(x) so its \[\frac{dy}{dx} = 2\sin(x) \times \cos(x) = 2\sin(x)\cos(x) = \sin(2x)\] hope it helps

OpenStudy (aum):

\[ \frac{d}{dx}\sin^2(x) = \frac{d}{dx}(\sin(x))^2 = 2(\sin(x))^{2-1} * \frac{d}{dx}(\sin(x) = 2\sin(x)*\cos(x) = \sin(2x) \]